2x^4-72x^2=0 what is the answer? this one is killing me.
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

2x^4-72x^2=0
----
Factor:
2x^2(x^2 - 36) = 0
Factor:
2x^2(x-6)(x+6) = 0
---
x = 0, x = 0 or x = 6 or x = -6
=============================
Cheers,
Stan H.

Answer by richwmiller(17219)

2x^4-72x^2=0
so what is the question? BTW that is one mighty equation? Do you let it out on weekends?
factor out 2x^2 then report back
it becomes difference of squares which can be factored